OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–2, –3) and is parallel to the line y = x + 9?

OpenStudy (anonymous):

slope of \(y=mx+b\) is \(m\) in your case \(m=1\) so your slope should be 1

OpenStudy (anonymous):

okayy , where do i go from there ?

OpenStudy (anonymous):

then use the point slope formula \(y-y_1=m(x-x_1)\) with \(x_1=-2, y_1=-3.m=1\)

OpenStudy (anonymous):

you should start with \[y+3=1(x+2)\] then get rid of the silly "1" and write \[y+3=x+2\]

OpenStudy (anonymous):

and finally you can solve for \(y\) by subtracting 3 from both sides and write \[y=x-1\]

OpenStudy (anonymous):

so then this would be my final equation ?